OpenStudy (anonymous):

is this true, because sin(-t)=-sin(t) it can be said that the sine of a negative angle is a negative number.

OpenStudy (anonymous):

Not always

OpenStudy (anonymous):

all sin(-t)=-sin(t) is saying is that sine is an odd function

OpenStudy (anonymous):

so its false?

OpenStudy (anonymous):

the angle -315 degrees is equivalent to positive 45 degrees

OpenStudy (kainui):

What is sin(-270)?

OpenStudy (anonymous):

and the sin of 45 is positive

OpenStudy (anonymous):

so it is false

OpenStudy (anonymous):

\(-1\le \sin \theta \le 1\) and the same will be true for \(-\theta\)